# Minimal Torsion Curves in Geometric Isogeny Classes

### v1

@@ 5.1. A Preliminary Result
Let $F=\mathbb{Q}(x)$. By Lemma 3.1, there is a model of $E/F$ where $P\in E(F)$ and such that there exists an $F$-rational cyclic isogeny $\varphi:E\rightarrow E^{\prime}$ with $j(E^{\prime})\in\mathbb{Q}$. Since $E$ has an $\ell$-isogeny over $F$, so does $E^{\prime}$ by [14, Proposition 3.2]. It follows from [14, Proposition 3.3] that any $E_{0}/\mathbb{Q}\in\mathcal{E}$ with $j(E_{0})=j(E^{\prime})$ has a $\mathbb{Q}$-rational cyclic $\ell$-isogeny, unless $\ell=7$ and $j(E_{0})=3^{3}\cdot 5\cdot 7^{5}/2^{7}$. Work of Mazur [31] implies $\ell\leq 37$. By applying Lemma 5.2 and checking the possible $x\in X_{1}(\ell)$ of odd degree associated to elliptic curves over $\mathbb{Q}$ in $\mathcal{E}$, where we may omit images as appearing in Remark 3.4, we see that we are done unless $\ell=7$ and $\mathcal{E}$ contains an elliptic curve with $j$-invariant $3^{3}\cdot 5\cdot 7^{5}/2^{7}$.
Now, it suffices to assume $\ell=7$ and $j(E_{0})=3^{3}\cdot 5\cdot 7^{5}/2^{7}$. Recall $P\in E(F)$ is a point of order $7^{k}$. Notice that $7^{k-1}P$ is a point of order $7$ on $E$ and is also defined over $F$ on $E$. By [8, Corollary 4.3], the curve $E^{\prime}$ has a rational point of order $7$ over an extension $F^{\prime}/F$ of degree $1$ or $7$. As $j(E^{\prime})=3^{3}\cdot 5\cdot 7^{5}/2^{7}$, a computation with division polynomials shows the residue field of a closed point on $X_{1}(7)$ associated to $E^{\prime}$ has degree 6 or 9, so $[F^{\prime}:\mathbb{Q}]$ is divisible by 6 or 9. Since $[F^{\prime}:F]$ divides 7, it follows that 6 or 9 must divide $[F:\mathbb{Q}]$. Since $F$ is an extension of odd degree, we must have $9\mid[F:\mathbb{Q}]$. Moreover, in [8, Proposition 4.1], it is proven that $3\cdot 7^{\text{max}(0,2k-3)}$ divides $[F:\mathbb{Q}]$. Therefore, $9\cdot 7^{\text{max}(0,2k-3)}$ divides $[F:\mathbb{Q}]$.
@@
Thus over a cubic extension $F$, the curve $E_{0}$ attains two independent 7 isogenies, and is $F$-isogenous to an elliptic curve $E_{1}/F$ with a rational cyclic 49-isogeny. By Proposition 2.3, the curve $E_{1}$ gives a closed point on $X_{1}(49)$ of degree dividing $9\cdot 7$. The previous paragraph shows it must have degree exactly $9\cdot 7$, and since $\deg(X_{1}(7^{k})\rightarrow X_{1}(7^{2}))$ has degree $7^{2k-4}$, the divisibility conditions of Proposition 3.2 are best-possible.
@@ 6. Results for non-CM classes and $\ell=3$
In this section, we will prove the following result, which includes instances where the divisibility conditions of Proposition 3.2 can be improved. It proves Theorem 1.1 if $\ell=3$.
@@
    Suppose there is no $E^{\prime}/\mathbb{Q}\in\mathcal{E}$ with $\operatorname{im}\rho_{E^{\prime},3^{\infty}}\in\{9.12.0.2,9.36.0.2,9.36.0.7,9.36.0.8\}$. For any $k\in\mathbb{Z}^{+}$, there exists a point of degree $\delta$ on $X_{1}(3^{k})$ associated to $j_{\text{min}}\in\mathcal{E}$.
@@
    Suppose there is $E^{\prime}/\mathbb{Q}\in\mathcal{E}$ with $\operatorname{im}\rho_{E^{\prime},3^{\infty}}\in\{9.12.0.2,9.36.0.7,9.36.0.8\}$. For $k=1$ or $k\geq 3$, there exists a point of degree $\delta$ on $X_{1}(3^{k})$ associated to $j_{\text{min}}\in\mathcal{E}$. If $k=2$, then $3\delta\mid\deg(x)$ and there exists a point of degree $3\delta$ on $X_{1}(3^{k})$ associated to $j_{\text{min}}\in\mathcal{E}$.
@@
    Suppose there is $E^{\prime}/\mathbb{Q}\in\mathcal{E}$ with $\operatorname{im}\rho_{E^{\prime},3^{\infty}}=9.36.0.2$. For $k=1$ or $k\geq 4$, there exists a point of degree $\delta$ on $X_{1}(3^{k})$ associated to $j_{\text{min}}\in\mathcal{E}$. Otherwise $3\delta\mid\deg(x)$ and there exists a point of degree $3\delta$ on $X_{1}(3^{k})$ associated to $j_{\text{min}}\in\mathcal{E}$.
One can take $j_{\text{min}}\in\mathbb{Q}$, unless we are in case (2) with $k>2$ or case (3) with $k>3$; in these latter cases one can take $j_{\text{min}}$ to be $3$-isogenous to a rational $j$-invariant.
@@
Let $\mathcal{E}$ be a rational $\overline{\mathbb{Q}}$-isogeny class of non-CM elliptic curves. If $E\in\mathcal{E}$ and $x=[E,P]\in X_{1}(3^{k})$ is a point of odd degree, then $3^{\max(0,2k-4)}\mid\deg(x)$, and this is best possible without placing restrictions on $\mathcal{E}$.
@@ 6.1. Preliminary Results
Suppose $3^{\max(0,2k-1-d)}\nmid[F:\mathbb{Q}]$, so in particular $02k-1-d>$ 0. Let $\{P,Q\}$ be a basis of $E[3^{k}]$. Replacing $F$ with at worst a quadratic extension $L/F$, we may view $\varphi$ as an $L$-isogeny from $E$ to $E_{0}/L$. Then $\operatorname{im}\rho_{E/L,3^{k}}$ is contained in
@@
By [8, Lemma 4.5], we have $3^{2k-1}\mid[\operatorname{GL}_{2}(\mathbb{Z}_{3}):\operatorname{im}\rho_{E_{0}/\mathbb{Q},3^{\infty}}]\cdot[L\cap\mathbb{Q}(E_{0}[3^{\infty}]):\mathbb{Q}]$. Since
@@
Note $\operatorname{im}\rho_{E/F,3^{k}}$ contained in $H$ as well, and since $L/F$ is at worst a quadratic extension, we have $\operatorname{ord}_{3}(\#\operatorname{im}\rho_{E/F,3^{k}})={2k-1}$. If $\operatorname{im}\rho_{E/F,3^{k}}$ is properly contained in $H$, then #$\#\operatorname{im}\rho_{E/F,3^{k}}=3^{2k-1}$. Since $\mathbb{Q}(\zeta_{3^{k}})\subseteq F$, it follows that 2 must divide $[F:\mathbb{Q}]$. This contradicts $F$ having odd degree. Hence $\operatorname{im}\rho_{E/F,3^{k}}=H$.
If $\operatorname{im}\rho_{E/F,3^{\infty}}\neq\pi^{-1}(\operatorname{im}\rho_{E/F,3^{k}})$, then $[F(E[3^{k+1}]):F(E[3^{k}])]$ divides $3^{3}$; see, for example, [6, Proposition 3.5]. Since #$\#\operatorname{Gal}(F(E[3^{k}])/F)=3^{2k-1}\cdot 2$, we have
#$\#\operatorname{Gal}(F(E[3^{k+1}])/F)\mid 3^{3}\cdot 3^{2k-1}\cdot 2=3^{2k+2}\cdot 2.$
It follows that #$\#\operatorname{Gal}(L(E[3^{k+1}])/L)\mid 3^{2k+2}\cdot 2$, and so the index of the 3-adic Galois representation of $E/L$ is divisible by at least $3^{2k-1}\cdot 8$. By [8, Lemma 4.5], we have $3^{2k-1}\cdot 8\mid[\operatorname{GL}_{2}(\mathbb{Z}_{3}):\operatorname{im}\rho_{E_{0}/\mathbb{Q},3^{\infty}}]\cdot[L\cap\mathbb{Q}(E_{0}[3^{\infty}]):\mathbb{Q}]$. It follows that $3^{2k-1-d}\mid[L\cap\mathbb{Q}(E_{0}[3^{\infty}]):\mathbb{Q}]$. Since $L$ is at worst a quadratic extension of $F$, then $3^{2k-1-d}\mid[F:\mathbb{Q}]$, contradicting our assumption. Thus $\operatorname{im}\rho_{E/F,3^{\infty}}=\pi^{-1}(\operatorname{im}\rho_{E/F,3^{k}})$
@@
Suppose $3^{\max(0,2k-1-d)}\nmid[F:\mathbb{Q}]$, so $02k-1-d>$ 0. First, suppose for the sake of contradiction that $r>k$. Then
@@
By [14, Lemma A.1], we may assume $\varphi$ is cyclic and generated by a point of order $3^{r}\cdot n$ where $3\nmid n$. If $n>1$, then by replacing $E$ with an $n$-isogenous curve if necessary, we may assume $\varphi$ has degree $3^{r}$. If $\operatorname{im}\rho_{E_{0},3^{\infty}}=9.36.0.6$ or $9.36.0.8$, then $\operatorname{ord}_{3}([\operatorname{GL}_{2}(\mathbb{Z}_{3}):\operatorname{im}\rho_{E_{0},3^{\infty}}])=2$. Suppose for the sake of contradiction that $3^{2k-3}\nmid[F:\mathbb{Q}]$. By Lemma 6.5, we have $r\leq k$ and $\ker(\varphi)\subseteq\langle P\rangle$.
First, suppose $r=k-1$ or $k$. Set $t=r+2$, and let $\{R,S\}$ be a basis of $E[3^{t}]$ such that $3^{t-k}R=P$ and $3^{t-k}S=Q$. Then we will show $\{\varphi(R),3^{k-2}\varphi(Q)\}$ is a basis of $E^{\prime}[9]$. Suppose there exist $\alpha,\beta\in\mathbb{Z}/9\mathbb{Z}$ such that $\alpha\varphi(R)=\beta 3^{k-2}\varphi(Q)$. This implies $\alpha R-\beta 3^{k-2}Q\in\ker(\varphi)\subseteq\langle P\rangle=\langle 3^{t-k}R\rangle$. Thus $\beta 3^{t-s}S$ is in the cyclic subgroup generated by $R$, and so $3^{2}\mid\beta$ since $\{R,S\}$ is a basis of $E[3^{t}]$. Thus $\alpha\varphi(R)=\beta 3^{k-2}\varphi(Q)=\mathcal{O}$, as desired.
@@
Let $F=\mathbb{Q}(x)$. By Lemma 3.1, there is a model of $E/F$ where $P\in E(F)$ and such that there exists an $F$-rational cyclic isogeny $\varphi:E\rightarrow E^{\prime}$ with $j(E^{\prime})\in\mathbb{Q}$. Replacing $E$ with an isogenous curve if necessary, we may assume $\varphi$ has degree $3^{r}$. Since $E$ has a 3-isogeny over $F$, so does $E^{\prime}$ by [14, Proposition 3.2]. It follows from [14, Proposition 3.3] that any $E_{0}/\mathbb{Q}$ with $j(E_{0})=j(E^{\prime})$ has a $\mathbb{Q}$-rational cyclic 3-isogeny, and $E_{0}$ gives a degree 1 closed point on $X_{1}(3)$ by Proposition 2.3. By [27, Corollary 1.3.1], the 3-adic image is one of the following groups, and we will consider each separately. Since we are interested in closed points on modular curves, and the degree of these points is not altered by taking quadratic twists, we may restrict to cases where $-I$ is contained in the 3-adic image.
@@
    $\operatorname{im}\rho_{E_{0},3^{\infty}}=3.4.0.1$: Since $d=0$ and $\deg(X_{1}(3^{k})\rightarrow X_{1}(3))=3^{2k-2}$, the divisibility condition of Proposition 3.2 is best possible for all $k\in\mathbb{Z}^{+}$, and one can take $j_{\text{min}}=j(E_{0})$.
@@
    $\operatorname{im}\rho_{E_{0},3^{\infty}}=3.12.0.1$ or $9.12.0.1$: If $\operatorname{im}\rho_{E_{0},3^{\infty}}=3.12.0.1$, then $E_{0}$ is 3-isogenous to an elliptic curve $E^{\prime}/\mathbb{Q}$ with a rational cyclic 9-isogeny. By Proposition 3.5, the 3-adic Galois representation of $E^{\prime}$ is completely determined by $\rho_{E_{0},3^{\infty}}$, and so it suffices to check a specific example. By viewing isogeny class 175.b in the LMFDB, we see that $\operatorname{im}\rho_{E^{\prime},3^{\infty}}=9.12.0.1$. Replacing $E_{0}$ with $E^{\prime}$ if necessary, we are free to assume $E_{0}$ has image 9.12.0.1. Thus $E_{0}$ corresponds to closed points on $X_{1}(3)$ and $X_{1}(9)$ of degree 1 and 3, respectively. Since $d=1$ and $\deg(X_{1}(3^{k})\rightarrow X_{1}(3))=3^{2k-2}$ for $k\geq 2$, the divisibility condition of Proposition 3.2 is best-possible for $k\in\mathbb{Z}^{+}$ and we can take $j_{\text{min}}=j(E_{0})$.
@@
    $\operatorname{im}\rho_{E_{0},3^{\infty}}=9.12.0.2$: A Magma computation shows that for $E_{0}/\mathbb{Q}$ with this image, there exists a cubic extension $L$ such that $E_{0}/L$ has an $L$-rational 9-isogeny and an independent 3-isogeny. Thus over $L$, the curve $E_{0}$ is 3-isogenous to $E_{1}/L$ with a rational cyclic 27-isogeny. Thus $E_{1}$ gives a closed point on $X_{1}(27)$ of degree at most 27 by Proposition 2.3. Since $d=1$ and $\deg(X_{1}(3^{k})\rightarrow X_{1}(27))=3^{2k-6}$, the divisibility conditions of Proposition 3.2 are best-possible for all $k\geq 3$ with $j_{\text{min}}=j(E_{1})$. No elliptic curve in $\mathcal{E}$ with $j\in\mathbb{Q}$ has a point of order 27 in this degree or lower.
@@
    $\operatorname{im}\rho_{E_{0},3^{\infty}}=9.36.0.2$ or $27.36.0.1$: As in case (2), the isogeny class 304.c in the LMFDB shows we are free to assume $E_{0}$ has image 27.36.0.1. A Magma computation shows that for $E_{0}/\mathbb{Q}$ with this image, there exists a cubic extension $L$ such that $E_{0}/L$ has an $L$-rational 27-isogeny and an independent 3-isogeny. Thus over $L$, the curve $E_{0}$ is 3-isogenous to $E_{1}/L$ with a rational cyclic 81-isogeny. Thus $E_{1}$ gives a closed point on $X_{1}(81)$ of degree at most 81 by Proposition 2.3. Since $d=2$ and $\deg(X_{1}(3^{k})\rightarrow X_{1}(81))=3^{2k-8}$, the divisibility conditions of Proposition 3.2 are best-possible for all $k\geq 4$ with $j_{\text{min}}=j(E_{1})$. No elliptic curve in $\mathcal{E}$ with $j$-invariant in $\mathbb{Q}$ has a point of order 81 in this degree or lower.
@@
    $\operatorname{im}\rho_{E_{0},3^{\infty}}=9.36.0.3$ or $9.36.0.6$: As in case (2), the isogeny class 22491.u in the LMFDB shows we are free to assume $E_{0}$ has image 9.36.0.6. Then $3^{\max(0,2k-3)}\mid[F:\mathbb{Q}]$ by Proposition 6.6. The conclusion follows with $j_{\text{min}}=j(E_{0})$.
@@
    $\operatorname{im}\rho_{E_{0},3^{\infty}}=9.36.0.1$, $9.36.0.4$, or $9.36.0.5$: As in case (2), the isogeny class 432.b in the LMFBD shows we are free to assume $E_{0}$ has image 9.36.0.4. A twist of $E_{0}$ has a rational point of order 9 and $d=2$, so divisibility conditions of Proposition 3.2 are best possible for all $k\in\mathbb{Z}^{+}$ with $j_{\text{min}}=j(E_{0})$.

### v2

@@ 5.1. A Preliminary Result
Let $F=\mathbb{Q}(x)$. By Lemma 3.1, there is a model of $E/F$ where $P\in E(F)$ and such that there exists an $F$-rational cyclic isogeny $\varphi:E\rightarrow E^{\prime}$ with $j(E^{\prime})\in\mathbb{Q}$. Since $E$ has an $\ell$-isogeny over $F$, so does $E^{\prime}$ by [CremonaNajmanQCurve, Proposition 3.2]. It follows from [CremonaNajmanQCurve, Proposition 3.3] that any $E_{0}/\mathbb{Q}\in\mathcal{E}$ with $j(E_{0})=j(E^{\prime})$ has a $\mathbb{Q}$-rational cyclic $\ell$-isogeny, unless $\ell=7$ and $j(E_{0})=3^{3}\cdot 5\cdot 7^{5}/2^{7}$. Work of Mazur [mazur78] implies $\ell\leq 37$. By applying Lemma 5.3 and checking the possible $x\in X_{1}(\ell)$ of odd degree associated to elliptic curves over $\mathbb{Q}$ in $\mathcal{E}$, where we may omit images as appearing in Remark 3.4, we see that we are done unless $\ell=7$ and $\mathcal{E}$ contains an elliptic curve with $j$-invariant $3^{3}\cdot 5\cdot 7^{5}/2^{7}$.
Let us discuss the case $\ell=7$ and $j(E_{0})=3^{3}\cdot 5\cdot 7^{5}/2^{7}$. Recall $P\in E(F)$ is a point of order $7^{k}$. Notice that $7^{k-1}P$ is a point of order $7$ on $E$ and is also defined over $F$ on $E$. By [OddDegQCurve, Corollary 4.3], the curve $E^{\prime}$ has a rational point of order $7$ over an extension $F^{\prime}/F$ of degree $1$ or $7$. As $j(E^{\prime})=3^{3}\cdot 5\cdot 7^{5}/2^{7}$, a computation with division polynomials shows the residue field of a closed point on $X_{1}(7)$ associated to $E^{\prime}$ has degree 6 or 9, so $F^{\prime}:\mathbb{Q}]$ is divisible by 6 or 9. Since $F^{\prime}:F]$ divides 7, it follows that 6 or 9 must divide $F:\mathbb{Q}]$. Since $F$ is an extension of odd degree, we must have $9\mid[F:\mathbb{Q}]$. Moreover, in [OddDegQCurve, Proposition 4.1], it is proven that $3\cdot 7^{\text{max}(0,2k-3)}$ divides $F:\mathbb{Q}]$. Therefore, $9\cdot 7^{\text{max}(0,2k-3)}$ divides $F:\mathbb{Q}]$.
@@
Thus over a cubic extension $F$, the curve $E_{0}$ attains two independent 7 isogenies, and is $F$-isogenous to an elliptic curve $E_{1}/F$ with a rational cyclic 49-isogeny. By Proposition 2.3, the curve $E_{1}$ gives a closed point on $X_{1}(49)$ of degree dividing $9\cdot 7$. The previous paragraph shows it must have degree exactly $9\cdot 7$, and since $\deg(X_{1}(7^{k})\rightarrow X_{1}(7^{2}))$ has degree $7^{2k-4}$, the divisibility conditions of Proposition 3.2 are the best-possible.
@@ 6. Results for non-CM classes and $\ell=3$
In this section, we will prove the following result, which includes instances where the divisibility conditions of Proposition 3.2 can be improved. It proves Theorem 1.2 if $\ell=3$. See Section 2.1 for a discussion of the notation used for subgroups of $\operatorname{GL}_{2}(\mathbb{Z}_{3})$.
@@
    Suppose there is no $E^{\prime}/\mathbb{Q}\in\mathcal{E}$ with $\operatorname{im}\rho_{E^{\prime},3^{\infty}}\in\{9.12.0.2,9.36.0.2,9.36.0.7,9.36.0.8\}$. For any $k\in\mathbb{Z}^{+}$, there exists a point of degree $\delta$ on $X_{1}(3^{k})$ associated to $j_{min}\in\mathcal{E}$.
@@
    Suppose there is an $E^{\prime}/\mathbb{Q}\in\mathcal{E}$ with $\operatorname{im}\rho_{E^{\prime},3^{\infty}}\in\{9.12.0.2,9.36.0.7,9.36.0.8\}$. For $k=1$ or $k\geq 3$, there exists a point of degree $\delta$ on $X_{1}(3^{k})$ associated to $j_{\min}\in\mathcal{E}$. If $k=2$, then $3\delta\mid\deg(x)$ and there exists a point of degree $3\delta$ on $X_{1}(3^{k})$ associated to $j_{min}\in\mathcal{E}$.
@@
    Suppose there is an $E^{\prime}/\mathbb{Q}\in\mathcal{E}$ with $\operatorname{im}\rho_{E^{\prime},3^{\infty}}=9.36.0.2$. For $k=1$ or $k\geq 4$, there exists a point of degree $\delta$ on $X_{1}(3^{k})$ associated to $j_{min}\in\mathcal{E}$. Otherwise $3\delta\mid\deg(x)$ and there exists a point of degree $3\delta$ on $X_{1}(3^{k})$ associated to $j_{min}\in\mathcal{E}$.
One can take $j_{min}\in\mathbb{Q}$, unless we are in case (2) with $k>2$ or case (3) with $k>3$; in these latter cases one can take $j_{min}$ to be $3$-isogenous to a rational $j$-invariant.
@@
Let $\mathcal{E}$ be a rational $\overline{\mathbb{Q}}$-isogeny class of non-CM elliptic curves. If $E\in\mathcal{E}$ and $x=[E,P]\in X_{1}(3^{k})$ is a point of odd degree, then $3^{\max(0,2k-4)}\mid\deg(x)$, and this is the best possible across all such $\mathcal{E}$.
@@ 6.1. Preliminary Results
Suppose $3^{\max(0,2k-1-d)}\nmid[F:\mathbb{Q}]$, so in particular 2k1d>02k-1-d>0. Let $\{P,Q\}$ be a basis of $E[3^{k}]$. Replacing $F$ with at worst a quadratic extension $L/F$, we may view $\varphi$ as an $L$-isogeny from $E$ to $E_{0}/L$. Then $\operatorname{im}\rho_{E/L,3^{k}}$ is contained in
@@
By [OddDegQCurve, Lemma 4.5], we have $3^{2k-1}\mid[\operatorname{GL}_{2}(\mathbb{Z}_{3}):\operatorname{im}\rho_{E_{0}/\mathbb{Q},3^{\infty}}]\cdot[L\cap\mathbb{Q}(E_{0}[3^{\infty}]):\mathbb{Q}]$. Since
@@
Note $\operatorname{im}\rho_{E/F,3^{k}}$ contained in $H$ as well, and since $L/F$ is at worst a quadratic extension, we have $\operatorname{ord}_{3}(\#\operatorname{im}\rho_{E/F,3^{k}})={2k-1}$. If $\operatorname{im}\rho_{E/F,3^{k}}$ is properly contained in $H$, then $\operatorname{im}\rho_{E/F,3^{k}}=3^{2k-1}$. Since $\mathbb{Q}(\zeta_{3^{k}})\subseteq F(E[3^{k}])$, it follows that 2 must divide $F:\mathbb{Q}]$. This contradicts $F$ having odd degree. Hence $\operatorname{im}\rho_{E/F,3^{k}}=H$.
If $\operatorname{im}\rho_{E/F,3^{\infty}}\neq\pi^{-1}(\operatorname{im}\rho_{E/F,3^{k}})$, then $F(E[3^{k+1}]):F(E[3^{k}])]$ divides $3^{3}$; see, for example, [BELOV, Proposition 3.5]. Since $\operatorname{Gal}(F(E[3^{k}])/F)=3^{2k-1}\cdot 2$, we have
$\operatorname{Gal}(F(E[3^{k+1}])/F)\mid 3^{3}\cdot 3^{2k-1}\cdot 2=3^{2k+2}\cdot 2.$
It follows that $\operatorname{Gal}(L(E[3^{k+1}])/L)\mid 3^{2k+2}\cdot 2$, and so the index of the 3-adic Galois representation of $E/L$ is divisible by at least $3^{2k-1}\cdot 8$. By [OddDegQCurve, Lemma 4.5], we have $3^{2k-1}\cdot 8\mid[\operatorname{GL}_{2}(\mathbb{Z}_{3}):\operatorname{im}\rho_{E_{0}/\mathbb{Q},3^{\infty}}]\cdot[L\cap\mathbb{Q}(E_{0}[3^{\infty}]):\mathbb{Q}]$. It follows that $3^{2k-1-d}\mid[L\cap\mathbb{Q}(E_{0}[3^{\infty}]):\mathbb{Q}]$. Since $L$ is at worst a quadratic extension of $F$, then $3^{2k-1-d}\mid[F:\mathbb{Q}]$, contradicting our assumption. Thus, $\operatorname{im}\rho_{E/F,3^{\infty}}=\pi^{-1}(\operatorname{im}\rho_{E/F,3^{k}})$. ∎
@@
Suppose $3^{\max(0,2k-1-d)}\nmid[F:\mathbb{Q}]$, so 2k1d>02k-1-d>0. First, suppose for the sake of contradiction that $r>k$. Then
@@
By [CremonaNajmanQCurve, Lemma A.1], we may assume $\varphi$ is cyclic and generated by a point of order $3^{r}\cdot n$ where $3\nmid n$. If $n>1$, then by replacing $E$ with an $n$-isogenous curve if necessary, we may assume $\varphi$ has degree $3^{r}$. If $\operatorname{im}\rho_{E_{0},3^{\infty}}=9.36.0.6$ or $9.36.0.8$, then $\operatorname{ord}_{3}([\operatorname{GL}_{2}(\mathbb{Z}_{3}):\operatorname{im}\rho_{E_{0},3^{\infty}}])=2$. Suppose for the sake of contradiction that $3^{2k-3}\nmid[F:\mathbb{Q}]$. By Lemma 6.5, we have $r\leq k$ and $\ker(\varphi)\subseteq\langle P\rangle$.
First, suppose r=k1r=k-1 or $k$. Set $t=r+2$, and let $\{R,S\}$ be a basis of $E[3^{t}]$ such that $3^{t-k}R=P$ and $3^{t-k}S=Q$. Then we will show $\{\varphi(R),3^{k-2}\varphi(Q)\}$ is a basis of $E^{\prime}[9]$. Suppose there exist $\alpha,\beta\in\mathbb{Z}/9\mathbb{Z}$ such that $\alpha\varphi(R)=\beta 3^{k-2}\varphi(Q)$. This implies $\alpha R-\beta 3^{k-2}Q\in\ker(\varphi)\subseteq\langle P\rangle=\langle 3^{t-k}R\rangle$. Thus $\beta 3^{t-s}S$ is in the cyclic subgroup generated by $R$, and so $3^{2}\mid\beta$ since $\{R,S\}$ is a basis of $E[3^{t}]$. Thus $\alpha\varphi(R)=\beta 3^{k-2}\varphi(Q)=\mathcal{O}$, as desired.
@@
Let $F=\mathbb{Q}(x)$. By Lemma 3.1, there is a model of $E/F$ where $P\in E(F)$ and such that there exists an $F$-rational cyclic isogeny $\varphi:E\rightarrow E^{\prime}$ with $j(E^{\prime})\in\mathbb{Q}$. Replacing $E$ with an isogenous curve if necessary, we may assume $\varphi$ has degree $3^{r}$. Since $E$ has a 3-isogeny over $F$, so does $E^{\prime}$ by [CremonaNajmanQCurve, Proposition 3.2]. It follows from [CremonaNajmanQCurve, Proposition 3.3] that any $E_{0}/\mathbb{Q}$ with $j(E_{0})=j(E^{\prime})$ has a $\mathbb{Q}$-rational cyclic 3-isogeny, and $E_{0}$ gives a degree 1 closed point on $X_{1}(3)$ by Proposition 2.3. By [RSZ21, Corollary 1.3.1], the 3-adic image is one of the following groups, and we will consider each separately. Since we are interested in closed points on modular curves, and the degree of these points is not altered by taking quadratic twists, we may restrict to cases where I-I is contained in the 3-adic image.
@@
    $\operatorname{im}\rho_{E_{0},3^{\infty}}=3.4.0.1$: Since $d=0$ and $\deg(X_{1}(3^{k})\rightarrow X_{1}(3))=3^{2k-2}$, the divisibility condition of Proposition 3.2 is best possible for all $k\in\mathbb{Z}^{+}$, and one can take $j_{min}=j(E_{0})$.
@@
    $\operatorname{im}\rho_{E_{0},3^{\infty}}=3.12.0.1$ or $9.12.0.1$: If $\operatorname{im}\rho_{E_{0},3^{\infty}}=3.12.0.1$, then $E_{0}$ is 3-isogenous to an elliptic curve $E^{\prime}/\mathbb{Q}$ with a rational cyclic 9-isogeny. By Proposition 3.5, the 3-adic Galois representation of $E^{\prime}$ is completely determined by $\rho_{E_{0},3^{\infty}}$, and so it suffices to check a specific example. By viewing isogeny class 175.b in the LMFDB, we see that $\operatorname{im}\rho_{E^{\prime},3^{\infty}}=9.12.0.1$. Replacing $E_{0}$ with $E^{\prime}$ if necessary, we are free to assume $E_{0}$ has image 9.12.0.1. Thus $E_{0}$ corresponds to closed points on $X_{1}(3)$ and $X_{1}(9)$ of degree 1 and 3, respectively. Since $d=1$ and $\deg(X_{1}(3^{k})\rightarrow X_{1}(3))=3^{2k-2}$ for $k\geq 2$, the divisibility condition of Proposition 3.2 is best-possible for $k\in\mathbb{Z}^{+}$ and we can take $j_{min}=j(E_{0})$.
@@
    $\operatorname{im}\rho_{E_{0},3^{\infty}}=9.12.0.2$: A Magma computation shows that for $E_{0}/\mathbb{Q}$ with this image, there exists a cubic extension $L$ such that $E_{0}/L$ has an $L$-rational 9-isogeny and an independent 3-isogeny. Thus over $L$, the curve $E_{0}$ is 3-isogenous to $E_{1}/L$ with a rational cyclic 27-isogeny. Thus $E_{1}$ gives a closed point on $X_{1}(27)$ of degree at most 27 by Proposition 2.3. Since $d=1$ and $\deg(X_{1}(3^{k})\rightarrow X_{1}(27))=3^{2k-6}$, the divisibility conditions of Proposition 3.2 are best-possible for all $k\geq 3$ with $j_{min}=j(E_{1})$. No elliptic curve in $\mathcal{E}$ with $j\in\mathbb{Q}$ has a point of order 27 in this degree or lower.
@@
    $\operatorname{im}\rho_{E_{0},3^{\infty}}=9.36.0.2$ or $27.36.0.1$: As in case (2), the isogeny class 304.c in the LMFDB shows we are free to assume $E_{0}$ has image 27.36.0.1. A Magma computation shows that for $E_{0}/\mathbb{Q}$ with this image, there exists a cubic extension $L$ such that $E_{0}/L$ has an $L$-rational 27-isogeny and an independent 3-isogeny. Thus over $L$, the curve $E_{0}$ is 3-isogenous to $E_{1}/L$ with a rational cyclic 81-isogeny. Thus $E_{1}$ gives a closed point on $X_{1}(81)$ of degree at most 81 by Proposition 2.3. Since $d=2$ and $\deg(X_{1}(3^{k})\rightarrow X_{1}(81))=3^{2k-8}$, the divisibility conditions of Proposition 3.2 are best-possible for all $k\geq 4$ with $j_{min}=j(E_{1})$. No elliptic curve in $\mathcal{E}$ with $j$-invariant in $\mathbb{Q}$ has a point of order 81 in this degree or lower.
@@
    $\operatorname{im}\rho_{E_{0},3^{\infty}}=9.36.0.3$ or $9.36.0.6$: As in case (2), the isogeny class 22491.u in the LMFDB shows we are free to assume $E_{0}$ has image 9.36.0.6. Then $3^{\max(0,2k-3)}\mid[F:\mathbb{Q}]$ by Proposition 6.6. The conclusion follows with $j_{min}=j(E_{0})$.
@@
    $\operatorname{im}\rho_{E_{0},3^{\infty}}=9.36.0.1$, $9.36.0.4$, or $9.36.0.5$: As in case (2), the isogeny class 432.b in the LMFBD shows we are free to assume $E_{0}$ has image 9.36.0.4. A twist of $E_{0}$ has a rational point of order 9 and $d=2$, so divisibility conditions of Proposition 3.2 are best possible for all $k\in\mathbb{Z}^{+}$ with $j_{min}=j(E_{0})$.